OpenStudy (kaynick):

The points (3,r) and (−7,3) lie on a line with slope 3. Find the missing coordinate r

OpenStudy (caozeyuan):

dy/dx=y2-y1/x2-x1

OpenStudy (caozeyuan):

so 3-r/-7-3=3

OpenStudy (anonymous):

general equation for a line is : y=ax+b a is the slope, hence a=3 We know that when x=-7, y=3 let's plug these values to the general equation: Y=3*x+b >>> 3= 3*(-7) +b therefore b is 24 know let's find r for x=3 r=3*3+24=33